# Neurosymbolic Grounding for Compositional World Models

### v1

@@ 5 Related Work
Object-Oriented World Models: Object-oriented world models [21, 44, 34, 9, 36] are constructed to learn structured latent spaces that facilitate efficient dynamics prediction in environments that can be decomposed into unique objects. We highlight two primary themes in this domain.
@@
Interaction Modeling and Modularity in Representations: [21] offers a contrastive world modeling technique using graph neural networks on an object-factored representation. However, GNNs may introduce compounding errors in sparse interaction domains. Addressing this, [10] introduces neural production systems (NPS) for modeling sparse interactions, akin to repeated dynamic instantiation GNN edge instantiation. Cosmos is influenced by the NPS architecture, with distinctions highlighted in Figure 2. In parallel, [4] studies a hierarchical abstraction for world models, decomposing slots into a (dynamics relevant) state vector and a (dynamics invariant) type vector, with dynamics prediction focusing on the state vector. Cosmos’s symbolic abstraction is a generalization of this abstraction. Instead of maintaining a single “type vector”, we maintain a set of learnable symbol vectors and select a relevant subset for each entity. This allows Cosmos to naturally discover global and local invariances, utilizing them to route latent encodings (akin to “state vectors”) to appropriate transition modules.
@@
Compositional Generation and Equivariance in World Models: Equivariant neural networks harness group symmetries for efficient learning [28, 6, 7, 37, 26]. In the context of CG, [35] investigates constructing equivariant neural networks within the MDP state-action space. [44] establishes a connection between homomorphic MDPs and compositional generalization, expressing CG as a symmetry group of equivariance to object permutations and developing a world model equivariant to object permutation (using action attention). Cosmos adopts this idea, but integrates a modular transition model that also respects permutation equivariance in slot order.
@@
Vision-grounded Neurosymbolic Learning: Prior work in neurosymbolic learning has demonstrated that symbolic grounding helps facilitate domain generalization and data efficiency [1, 2, 30, 43, 29, 24, 16, 42]. The interplay between neural and symbolic encodings in these works can be abstracted into three categories: (1) scaffolding perceptual inputs with precomputed symbolic information for downstream prediction [24, 8, 2, 16, 33], (2) learning a symbolic abstraction from a perceptual input useful for downstream prediction [32], and (3) jointly learning a neural and symbolic encodings leveraged for prediction [43]. Our approach aligns most closely with the third category. While [43] combine neural and symbolic encoders in a VAE setting, highlighting the regularization advantages of symbols for unsupervised clustering, they rely on a program synthesizer to search for symbolic transformations in a DSL—introducing scalability and expressiveness issues. Cosmos also crafts a neurosymbolic encoding, but addresses scalability and expressiveness concerns of program synthesis by using a foundation model to generate the symbolic encodings.
@@
Foundation Models as Symbol Extractors: Many works employ foundation models to decompose complex visual reasoning tasks [38, 11, 25]. [16, 39] decompose natural language instructions into executable programs using a Code LLM for robotic manipulation and 3D understanding. Notably, [31] uses Code LLMs to decomposes natural language queries to API calls in a library of pretrained models. Such approaches necessitate hand-engineering the API to be expressive enough to generalize to all attributes in a domain. Cosmos builds upon the idea of using compositionality of symbols to execute parameterized modules but sidesteps the symbolic decomposition bottleneck by parsimoniously using the symbolic encodings only for selecting representative encodings. I.e., Cosmos does not need its symbols to learn perfect reconstructions. The symbolic encoding is only used for selecting modules, while the neural encoding can learn fine-grained dynamics-relevant attributes that may not be known ahead of time. Furthermore, to the best of our knowledge, Cosmos is the first work to leverage vision-language foundation models for compositional world modeling.
@@ 6 Conclusion
We have presented Cosmos, a new neurosymbolic approach for compositionally generalizable world modeling. Our two key findings are that annotating entities with symbolic attributes can help with CG and that it is possible to get these symbols “for free” from foundation models. We have considered two definitions of CG — one new to this paper — and show that: (i) CG world modeling still has a long way to go regarding performance, and (ii) neurosymbolic grounding helps enhance CG.
@@ 7 Acknowledgements
Research supported by The NSF National AI Institute for Foundations of Machine Learning (IFML), ARO award #W911NF-21-1-0009, and DARPA award #HR00112320018.
@@ Appendix A Appendix
The Appendix is divided into five sections. Section A.1 surveys the types of compositions we study in detail, Section A.2 introduces a faster algorithm used in implementation for module selection, Section A.3 outlines experiment details and, notably, presents justification for the Equivariant MRR metric employed to study encoding separation, Section A.4 presents details of how the downstream planning experiments were conducted, and Section A.5 showcases qualitative results on a randomly sampled subset of five-object state-action pairs.
A.1 Types of Compositions
@@
Figure 5: Overview of types of compositions studied. Scene composition (left) necessitates learning a world model that is equivariant to object replacement. Object compositions (right) necessitates learning the properties of scene composition as well as additional constraints where objects with shared attributes also have shared dynamics. We study two instantiations of shared attributes sets: “Sticky” and “Team”. Details on these instantiations are given in Appendix A.1.
@@
Scene Composition: Scene composition (Figure 5) necessitates learning a world model that is equivariant to object replacement. The dynamics of the environment depends on which objects are present in the scene.
@@
Object Composition: Object composition (Figure 5) necessitates learning all the properties present in scene composition. Additionally, in object composition, the composition is determined by constraints placed on observable attributes of individual objects. For instance, in Sticky block pushing (Fig 5), the scene is constrained so that two objects start out with the same color adjacent to each other; and an action on one object moves all objects of the same color with it. This gives the appearance of two objects being stuck to each other. At test time, the objects stuck together change. Sticky block pushing demonstrates compositionality constraints based on two attributes: position and color. In the team block pushing (Figure 5), we relax the adjacency constraint in the sticky block pushing domain. An action on any object also moves other objects of the same color. This allows us to study whether the adjacency constraint places a larger burden on dynamics learning than the color constraint.
A.2 Transition Algorithm
@@
NPS [9] necessitates selecting a primary slot $(p)$ to be modified, a contextual slot $(c)$, and a rule $(r)$ to modify the primary slot in the presence of the contextual slot. The naive algorithm to compute this tuple has a runtime of $O(k^{2}l)$ where $k$ is the number of slots and $l$ is the number of rules. However, in implementation, the selection of $(r,p,c)$ can be reduced to a runtime of $O(kl+k)$ by partial application of the query-key attention. This is achieved by selecting the primary slot $p$ and $\textsc{Mlp}_{i}$, partially transforming $S_{p}$ using a partial transition module $\textsc{Mlp}_{(i,\texttt{left})}$, selecting the contextual slot $c$, and performing a final transformation of $\textsc{Mlp}_{(i,\text{left})}(S_{p})$ with $S_{c}$ using $\textsc{Mlp}_{(i,\text{right})}$. Algorithm 2 presents this faster algorithm.
@@
1:function Transition(Key=$\{\overline{\Lambda}_{1},\overline{\Lambda}_{2},\overline{\Lambda}_{k}\}$, Query=$\{\vec{R}_{1},\vec{R}_{2},\vec{R}_{l}\}$, Value=$\{S_{1},S_{2},S_{k}\}$)
2:   $\textbf{A}^{\star}\leftarrow\textbf{GumbelSoftmax}(\textbf{KQAttention}(\text{key=}\{\overline{\Lambda}_{1},\overline{\Lambda}_{2},\overline{\Lambda}_{k}\},\text{query=}\{\vec{R}_{1},\vec{R}_{2},\vec{R}_{l}\}))$
@@
4:   $S^{\star}\leftarrow\textbf{LeftMLPBank}[r](\text{concat}(S_{p},\vec{R}_{r}))$
5:   $\textbf{A}_{2}^{\star}\leftarrow\textbf{GumbelSoftmax}(\textbf{KQAttention}(\text{key=}\{\overline{\Lambda}_{1},\overline{\Lambda}_{2},\overline{\Lambda}_{k}\},\text{query=}\{S^{\star},S^{\star},S^{\star}\}))$
@@
7:   $S^{\star}_{2}\leftarrow\textbf{RightMLPBank}[r](\text{concat}(S_{c},S^{\star}))$
@@
A.3 Evaluation Procedure
Dataset Generation: To generate each dataset, we first create a scene configuration file that prescribes the permissible shapes and colors for objects within a given dataset. The scene configuration ensures that $\mathcal{F}_{\mathtt{C}}(\mathcal{D}_{train})\cap\mathcal{F}_{\mathtt{C}}(\mathcal{D}_{eval})=\emptyset$. Next, we sample $\mathcal{D}_{train}$ and $\mathcal{D}_{eval}$ from the permissible scenes. Each dataset is a trajectory of state-action pairs, where the state is the image of the shape $3\times 224\times 224$ and the action is a vector, factorized by object ID (Object x North-East-South-West). Overall, we generate $3000$ trajectories of length $32$ each where the actions are sampled from a random uniform distribution. Our domain is equivalent to the observed weighted shapes setup studied in [19] with a compositionality constraint where the weight of the objects depend only on the shapes.
@@ A.3 Evaluation Procedure
Baselines: We compare against past works in compositional world modeling with publically available codebases at the time of writing. For the block pushing domain, we compare against Homomorphic world models (Gnn) [44] and an ablation of our model without symbols (AlignedNPS). HOWM uses a slot autoencoder, an action attention module and a GNN for modeling transitions. It requires a two-step training process: first warm starting the slot-autoencoder and then training the action attention model and GNN with an equivariant contrastive loss (Hungarian matching loss). NPS uses a slot autoencoder for modeling perceptions and a NPS module for modeling transitions. The pipeline is trained end to end with contrastive loss. We use action attention with NPS as well. For both models, we weren’t able to reproduce the results using the provided codebases due to issues in training robust perception models for large images ($3\times 224\times 224$). To ensure a fair comparison, and since both these methods are agnostic to the perception model, we opt to reimplement the core ideas for both these models and use the same fine-tuned perception model for all models.
@@
3 objects 5 objects Dataset Model MSE $\downarrow$ AE-MSE $\downarrow$ Eq.MRR $\uparrow$ MSE $\downarrow$ AE-MSE $\downarrow$ Eq.MRR $\uparrow$ OC Cosmos 4.23E-03 +/- 1.49E-04 4.90E-04 +/- 1.03E-04 1.20E-01 +/- 2.05E-02 4.15E-03 +/- 3.21E-03 1.68E-03 +/- 1.48E-03 3.67E-01 +/- 7.73E-02 (Sticky) AlignedNPS 1.14E-02 +/- 9.89E-04 7.72E-03 +/- 1.21E-03 8.01E-02 +/- 6.79E-02 6.07E-03 +/- 8.30E-04 2.47E-03 +/- 3.60E-04 3.62E-01 +/- 1.81E-02 Gnn 7.94E-03 +/- 5.47E-03 5.11E-03 +/- 4.94E-03 6.03E-04 +/- 1.02E-04 6.21E-03 +/- 1.26E-03 2.73E-03 +/- 1.27E-03 5.30E-04 +/- 5.15E-05 OC Cosmos 4.60E-03 +/- 2.32E-03 4.33E-04 +/- 1.58E-04 1.04E-01 +/- 3.19E-02 5.53E-03 +/- 1.95E-03 1.86E-03 +/- 1.61E-03 2.86E-01 +/- 4.32E-02 (Team) AlignedNPS 1.24E-02 +/- 4.11E-04 8.36E-03 +/- 6.78E-04 1.75E-01 +/- 2.68E-02 9.64E-03 +/- 1.95E-04 3.12E-03 +/- 6.07E-04 2.93E-01 +/- 2.02E-02 Gnn 8.92E-03 +/- 6.05E-03 3.82E-03 +/- 3.64E-03 7.16E-04 +/- 1.09E-04 7.01E-03 +/- 9.81E-04 1.62E-03 +/- 1.04E-03 5.46E-04 +/- 1.37E-04 SC Cosmos 7.66E-04 +/- 4.08E-04 6.34E-05 +/- 2.01E-05 2.99E-01 +/- 2.85E-02 4.08E-04 +/- 4.68E-06 2.92E-06 +/- 6.34E-07 3.03E-01 +/- 3.88E-02 AlignedNPS 3.51E-03 +/- 6.30E-04 2.69E-03 +/- 6.89E-05 2.97E-01 +/- 7.99E-02 2.45E-03 +/- 3.47E-04 1.22E-03 +/- 9.06E-04 3.19E-01 +/- 1.01E-01 Gnn 9.89E-03 +/- 5.77E-03 1.03E-02 +/- 5.44E-03 5.50E-01 +/- 5.18E-01 1.20E-02 +/- 1.13E-02 1.28E-02 +/- 1.08E-02 5.25E-01 +/- 2.67E-01
@@
Table 2: Evaluation results on the 2D block pushing domain for scene composition (SC) and object composition (OC) averaged across three seeds. This table includes standard deviation numbers as well. Our model (Cosmos) achieves best next-state reconstructions for all datasets.
A.4 Downstream Evaluation Setup

### v2

@@ 5 Related Work
Object-Oriented World Models: Object-oriented world models (Kipf et al., 2019; Zhao et al., 2022; Van der Pol et al., 2020a; Goyal et al., 2021; Veerapaneni et al., 2020) are constructed to learn structured latent spaces that facilitate efficient dynamics prediction in environments that can be decomposed into unique objects. We highlight two primary themes in this domain.
@@
Interaction Modeling and Modularity in Representations: Kipf et al. (2019) builds a contrastive world modeling technique using graph neural networks with an object-factored representation. However, GNNs may introduce compounding errors in sparse interaction domains. Addressing this, Goyal et al. (2021) introduces neural production systems (NPS) for modeling sparse interactions, akin to repeated dynamic instantiation GNN edge instantiation. Cosmos is influenced by the NPS architecture, with distinctions highlighted in Figure 2. In parallel, Chang et al. (2023) studies a hierarchical abstraction for world models, decomposing slots into a (dynamics relevant) state vector and a (dynamics invariant) type vector, with dynamics prediction focusing on the state vector. In Cosmos, instead of maintaining a single “type vector”, we maintain a set of learnable symbol vectors and select a relevant subset for each entity. This allows Cosmos to naturally discover global and local invariances, utilizing them to route latent encodings (akin to “state vectors”) to appropriate transition modules.
@@
Compositional Generalization and Equivariance in World Models: Equivariant neural networks harness group symmetries for efficient learning (Ravindran, 2004; Cohen & Welling, 2016a; b; Walters et al., 2020; Park et al., 2022). In the context of CompGen, Van der Pol et al. (2020b) investigates constructing equivariant neural networks within the MDP state-action space. Zhao et al. (2022) establishes a connection between homomorphic MDPs and compositional generalization, expressing CompGen as a symmetry group of equivariance to object permutations and developing a world model equivariant to object permutation (using action attention). We adopt this idea, but integrate a modular transition model that also respects permutation equivariance in slot order.
@@
Vision-grounded Neurosymbolic Learning: Prior work in neurosymbolic learning has demonstrated that symbolic grounding helps facilitate domain generalization and data efficiency (Andreas et al., 2016b; a; He et al., 2016; Zhan et al., 2021; Shah et al., 2020; Mao et al., 2019; Hsu et al., 2023; Yi et al., 2018). The interplay between neural and symbolic encodings in these works can be abstracted into three categories: (1) scaffolding perceptual inputs with precomputed symbolic information for downstream prediction (Mao et al., 2019; Ellis et al., 2018; Andreas et al., 2016a; Hsu et al., 2023; Valkov et al., 2018), (2) learning a symbolic abstraction from a perceptual input useful for downstream prediction Tang & Ellis (2023), and (3) jointly learning a neural and symbolic encodings leveraged for prediction Zhan et al. (2021). Our approach aligns most closely with the third category. While Zhan et al. (2021) combine neural and symbolic encoders in a VAE setting, highlighting the regularization advantages of symbols for unsupervised clustering, they rely on a program synthesizer to search for symbolic transformations in a DSL—introducing scalability and expressiveness issues. Cosmos also crafts a neurosymbolic encoding, but addresses scalability and expressiveness concerns of program synthesis by using a foundation model to generate the symbolic encodings.
@@
Foundation Models as Symbol Extractors: Many works employ foundation models to decompose complex visual reasoning tasks (Wang et al., 2023a; Gupta & Kembhavi, 2023; Nayak et al., 2022). (Hsu et al., 2023; Wang et al., 2023b) decompose natural language instructions into executable programs using a Code LLM for robotic manipulation and 3D understanding. Notably, ViperGPT (Surís et al., 2023) uses Code LLMs to decomposes natural language queries to API calls in a library of pretrained models. Such approaches necessitate hand-engineering the API to be expressive enough to generalize to all attributes in a domain. Cosmos builds upon the idea of using compositionality of symbols to execute parameterized modules but sidesteps the symbolic decomposition bottleneck by parsimoniously using the symbolic encodings only for selecting representative encodings. I.e., Cosmos does not need its symbols to learn perfect reconstructions. The symbolic encoding is only used for selecting modules, while the neural encoding can learn fine-grained dynamics-relevant attributes that may not be known ahead of time. Furthermore, to the best of our knowledge, Cosmos is the first work to leverage vision-language foundation models for compositional world modeling.
@@ 6 Conclusion
We have presented Cosmos, a new neurosymbolic approach for compositionally generalizable world modeling. Our two key findings are that annotating entities with symbolic attributes can help with CompGen and that it is possible to get these symbols “for free” from foundation models. We have considered two definitions of CompGen — one new to this paper — and show that: (i) CompGen world modeling still has a long way to go regarding performance, and (ii) neurosymbolic grounding helps enhance CompGen.
@@
7 Reproducibility and Ethical Considerations
7.1 Reproducibility Considerations
We encourage reproducibility of our work through the following steps:
Providing Code and Setup Environments: Upon acceptance of our work, we will release the complete source code, pretrained models, and associated setup environments under the MIT license. This is to ensure researchers can faithfully replicate our results without any hindrance.
Improving Baseline Reproducibility: Our methodology extensively utilizes these public repositories (Zhao et al., 2022; Kipf et al., 2019; Goyal et al., 2021) for generating data and computing evaluation metrics. During development of our method, we made several improvements to these repositories: (1) We enhanced the performance and usability of the models, (2) We updated the algorithms to accommodate the latest versions of the machine learning libraries, and (3) we constructed reproducible anaconda environments for each package to work on Ubuntu 22.04. We intend to contribute back by sending pull requests to the repositories with our improvements, including other bug fixes, runtime enhancements, and updates.
Outlining Methodology Details: The details of our model are meticulously described in Section 3. A comprehensive overview is available in Algorithm 1, and a deeper dive into the transition function is in Appendix Section A.3. We mention training details in appendix section A.4.
7.2 Ethical Considerations
Potential for Misuse: As with other ML techniques, world models can be harnessed by malicious actors to inflict societal harm. Our models are trained on synthetic block pushing datasets, which mitigates their potential for misuse.
Privacy Concerns: In the long term, as world models operate on unsupervised data, they enable learning behavioral profiles without the active knowledge or explicit consent of the subjects. This raises privacy issues, especially when considering real-world, non-synthetic datasets. We did not collect or retain any human-centric data during the course of this project.
Bias and Fairness: World models, generally, enable learning unbiased representations of data. However, we leverage foundation models which could be trained on biased data and such biases can reflect in our world models.
@@ 8 Acknowledgements
This research was supported by the NSF National AI Institute for Foundations of Machine Learning (IFML), ARO award #W911NF-21-1-0009, and DARPA award #HR00112320018.
@@ Appendix A Appendix
The Appendix is divided into eight sections. Section A.1 explains how we leverage SAM to generate a set-structured representation, Section A.2 surveys the types of compositions we study in detail, Section A.3 introduces a faster algorithm used in implementation for module selection, Section A.4 outlines experiment details and, notably, presents justification for the Equivariant MRR metric employed to study encoding separation, Section A.5 presents details of how the downstream planning experiments were conducted, Section A.6 goes over our reasoning for selecting relevant benchmarks, Section A.7 details an ablation study with a “fully-symbolic” model, and Section A.8 showcases qualitative results on a randomly sampled subset of five-object state-action pairs.
A.1 Generating set-structured representations with Segment Anything
We prompt SAM (Kirillov et al., 2023) with the image ($I$) and a $8\times 8$ grid of points. This yields $64\times 3$ potential masks (as there are three channels). To ensure a set structured representation, we must ensure that (1) each mask captures a specific property of the image, (2) collectively, all masks describe the entire image. We ensure (1) by removing duplicate masks and (2) by evaluating all combinations of remaining slots and selecting the $k$ tuple (where $k$ is the number of slots) that, when summed, most closely matches the image. The resulting masks $\{M_{1}^{\prime},\dots M_{k}^{\prime}\}$ are point-wise multiplied with the image to yield $\{I_{1},\dots I_{k}\}$. Each masked image is passed through a finetuned Resnet to yield $\{S_{1},\dots S_{k}\}$.
A.2 Types of Compositions
@@
Figure 5: Overview of types of compositions studied. Entity composition (left) necessitates learning a world model that is equivariant to object replacement. Relational compositions (right) necessitates learning the properties of entity composition as well as additional constraints where objects with shared attributes also have shared dynamics. We study two instantiations of shared attributes sets: “Sticky” and “Team”. Details on these instantiations are given in Appendix A.2.
@@
Entity Composition: Entity composition (Figure 5) necessitates learning a world model that is equivariant to object replacement. The dynamics of the environment depends on which objects are present in the scene.
@@
Relational Composition: Relational composition (Figure 5) necessitates learning all the properties present in entity composition. Additionally, in relational composition, the composition is determined by constraints placed on observable attributes of individual objects. For instance, in Sticky block pushing (Fig 5), the scene is constrained so that two objects start out with the same color adjacent to each other; and an action on one object moves all objects of the same color with it. This gives the appearance of two objects being stuck to each other. At test time, the objects stuck together change. Sticky block pushing demonstrates compositionality constraints based on two attributes: position and color. In the team block pushing (Figure 5), we relax the adjacency constraint in the sticky block pushing domain. An action on any object also moves other objects of the same color. This allows us to study whether the adjacency constraint places a larger burden on dynamics learning than the color constraint.
A.3 Transition Algorithm
@@
NPS (Goyal et al., 2021) necessitates selecting a primary slot $(p)$ to be modified, a contextual slot $(c)$, and a rule $(r)$ to modify the primary slot in the presence of the contextual slot. The naive algorithm to compute this tuple has a runtime of $O(k^{2}l)$ where $k$ is the number of slots and $l$ is the number of rules. However, in implementation, the selection of $(r,p,c)$ can be reduced to a runtime of $O(kl+k)$ by partial application of the query-key attention. This is achieved by selecting the primary slot $p$ and $\textsc{Mlp}_{i}$, partially transforming $S_{p}$ using a partial transition module $\textsc{Mlp}_{(i,\texttt{left})}$, selecting the contextual slot $c$, and performing a final transformation of $\textsc{Mlp}_{(i,\texttt{left})}(S_{p})$ with $S_{c}$ using $\textsc{Mlp}_{(i,\texttt{right})}$. Algorithm 2 presents this faster algorithm.
@@
1:function Transition(Key=$\{\overline{\Lambda}_{1},\dots\overline{\Lambda}_{k}\}$, Query=$\{\vec{R}_{1},\dots\vec{R}_{l}\}$, Value=$\{S_{1},\dots S_{k}\}$)
2:   $\textbf{A}^{\star}\leftarrow\textbf{GumbelSoftmax}(\textbf{KQAttention}(\text{key=}\{\overline{\Lambda}_{1},\dots\overline{\Lambda}_{k}\},\text{query=}\{\vec{R}_{1},\dots\vec{R}_{l}\}))$
@@
4:   $S^{\star}\leftarrow\textsc{Mlp}_{r,\texttt{left}}(\text{concat}(S_{p},\vec{R}_{r}))$
5:   $\textbf{A}_{2}^{\star}\leftarrow\textbf{GumbelSoftmax}(\textbf{KQAttention}(\text{key=}\{\overline{\Lambda}_{1},\dots\overline{\Lambda}_{k}\},\text{query=}\{S^{\star},S^{\star},S^{\star}\}))$
@@
7:   $S^{\star}_{2}\leftarrow\textsc{Mlp}_{r,\texttt{right}}(\text{concat}(S_{c},S^{\star}))$
@@
A.4 Evaluation Procedure
@@
Dataset Generation: To generate each dataset, we first create a scene configuration file that prescribes the permissible shapes and colors for objects within a given dataset. The scene configuration ensures that $\mathcal{F}_{\mathtt{C}}(\mathcal{D}_{train})\cap\mathcal{F}_{\mathtt{C}}(\mathcal{D}_{eval})=\emptyset$. Next, we sample $\mathcal{D}_{train}$ and $\mathcal{D}_{eval}$ from the permissible scenes. Each dataset is a trajectory of state-action pairs, where the state is the image of the shape $3\times 224\times 224$ and the action is a vector, factorized by object ID (Object x North-East-South-West). Overall, we generate $3000$ trajectories of length $32$ each where the actions are sampled from a random uniform distribution. Our domain is equivalent to the observed weighted shapes setup studied in Ke et al. (2021) with a compositionality constraint where the weight of the objects depend only on the shapes.
Baselines: We compare against past works in compositional world modeling with publically available codebases at the time of writing. For the block pushing domain, we compare against homomorphic world models (Gnn) (Zhao et al., 2022) and an ablation of our model without symbols (AlignedNPS). Gnn uses a slot autoencoder, an action attention module and a graph neural network for modeling transitions. It requires a two-step training process: first warm starting the slot-autoencoder and then training the action attention model and GNN with an equivariant contrastive loss (Hungarian matching loss). AlignedNPS uses a slot autoencoder for modeling perceptions and a NPS module (Goyal et al., 2021) for modeling transitions. The pipeline is trained end to end with contrastive loss. We use action attention with NPS as well. For both models, we weren’t able to reproduce the results using the provided codebases due to issues in training robust perception models for large images ($3\times 224\times 224$). To ensure a fair comparison, and since both these methods are agnostic to the perception model, we opt to reimplement the core ideas for both these models and use the same fine-tuned perception model for all models.
@@
3 objects 5 objects Dataset Model MSE $\downarrow$ AE-MSE $\downarrow$ Eq.MRR $\uparrow$ MSE $\downarrow$ AE-MSE $\downarrow$ Eq.MRR $\uparrow$ RC Cosmos 4.23E-03 +/- 1.49E-04 4.90E-04 +/- 1.03E-04 1.20E-01 +/- 2.05E-02 4.15E-03 +/- 3.21E-03 1.68E-03 +/- 1.48E-03 3.67E-01 +/- 7.73E-02 (Sticky) AlignedNPS 1.14E-02 +/- 9.89E-04 7.72E-03 +/- 1.21E-03 8.01E-02 +/- 6.79E-02 6.07E-03 +/- 8.30E-04 2.47E-03 +/- 3.60E-04 3.62E-01 +/- 1.81E-02 Gnn 7.94E-03 +/- 5.47E-03 5.11E-03 +/- 4.94E-03 6.03E-04 +/- 1.02E-04 6.21E-03 +/- 1.26E-03 2.73E-03 +/- 1.27E-03 5.30E-04 +/- 5.15E-05 RC Cosmos 4.60E-03 +/- 2.32E-03 4.33E-04 +/- 1.58E-04 1.04E-01 +/- 3.19E-02 5.53E-03 +/- 1.95E-03 1.86E-03 +/- 1.61E-03 2.86E-01 +/- 4.32E-02 (Team) AlignedNPS 1.24E-02 +/- 4.11E-04 8.36E-03 +/- 6.78E-04 1.75E-01 +/- 2.68E-02 9.64E-03 +/- 1.95E-04 3.12E-03 +/- 6.07E-04 2.93E-01 +/- 2.02E-02 Gnn 8.92E-03 +/- 6.05E-03 3.82E-03 +/- 3.64E-03 7.16E-04 +/- 1.09E-04 7.01E-03 +/- 9.81E-04 1.62E-03 +/- 1.04E-03 5.46E-04 +/- 1.37E-04 EC Cosmos 7.66E-04 +/- 4.08E-04 6.34E-05 +/- 2.01E-05 2.99E-01 +/- 2.85E-02 4.08E-04 +/- 4.68E-06 2.92E-06 +/- 6.34E-07 3.03E-01 +/- 3.88E-02 AlignedNPS 3.51E-03 +/- 6.30E-04 2.69E-03 +/- 6.89E-05 2.97E-01 +/- 7.99E-02 2.45E-03 +/- 3.47E-04 1.22E-03 +/- 9.06E-04 3.19E-01 +/- 1.01E-01 Gnn 9.89E-03 +/- 5.77E-03 1.03E-02 +/- 5.44E-03 5.50E-01 +/- 5.18E-01 1.20E-02 +/- 1.13E-02 1.28E-02 +/- 1.08E-02 5.25E-01 +/- 2.67E-01
@@
Table 2: Evaluation results on the 2D block pushing domain for entity composition (EC) and relational composition (RC) averaged across three seeds. This table includes standard deviation numbers as well. Our model (Cosmos) achieves best next-state reconstructions for all datasets.
A.5 Downstream Evaluation Setup